OpenStudy (anonymous):

2. Use the Law of sines to find the missing side of the triangle. Find the measure of b, given m

OpenStudy (anonymous):

@IMStuck

OpenStudy (johnweldon1993):

The law of sine states \[\large \frac{\sin(A)}{a} = \frac{\sin(B)}{b} = \frac{\sin(C)}{c}\] where A B and C are the angles given..and a b and c are the side lengths

OpenStudy (campbell_st):

well the law of sine is \[\frac{a}{\sin(A)} = \frac{b}{\sin(B)}\] so substitute your information and then solve for a hope it helps

OpenStudy (johnweldon1993):

if we know \[\large \frac{\sin(A)}{a} = \frac{\sin(38)}{31}\] and \[\large \frac{\sin(B)}{b} = \frac{\sin(74)}{b}\] we then know \[\large \frac{\sin(38)}{31} = \frac{\sin(74)}{b}\] and we can solve for 'b'

OpenStudy (anonymous):

18.3?

OpenStudy (johnweldon1993):

Hmm..not quite what I get... Rearranging that equation we have \[\large b = \frac{31\sin(74)}{\sin(38)} = ?\]

OpenStudy (anonymous):

now my calc says 37.8

OpenStudy (johnweldon1993):

Hmm still not what I'm getting..I receive an answer of \[\large 48.4\]

OpenStudy (anonymous):

Oh that's weird

OpenStudy (anonymous):

okay thanks!

OpenStudy (johnweldon1993):

No problem :)

OpenStudy (imstuck):

Is your calculator in the wrong mode!? If it is in radians and it needs to be in degrees, your answer will be WAYYYYY off!!! Check that and try again

OpenStudy (johnweldon1993):

My calculator?